OpenStudy (anonymous):

The sun produces 3.9 ⋅ 1033 ergs of radiant energy per second. How many ergs of radiant energy does the sun produce in 2.45 ⋅ 105 seconds?

OpenStudy (denonakavro):

Multiply all the numbers and tell me the answer

OpenStudy (whpalmer4):

This is a rate problem. You have a rate: \(3.9*10^{33} \text{ ergs/s}\) and you want to know the total produced if that rate is applied for a certain amount of time: \[x = r t\]where \(x\) is the amount produced, \(r\) is the rate and \(t\) is the time \[r = 3.9*10^{33}\text{ ergs/s}\]\[t=2.45*10^5\text{ s}\]\[x = r t = (3.9*10^{33}\text{ ergs/s})(2.45*10^5\text{ s})\]\[=(3.9*10^{33}\text{ ergs/}\cancel{\text{s}})(2.45*10^5\cancel{\text{ s}})\]\[=(3.9*10^{33})(2.45*10^5)\text{ ergs}\] Can you do that multiplication?